OpenStudy (jessiegonzales):

MEDAL!! ALGEBRA 2 19. l-4b-8l+l-1-b^2l +2b^3;b=2

OpenStudy (butterflydreamer):

\[\left| -4(2) - 8 \right| + \left| -1-(2)^2 \right| + 2(2)^3 \]

OpenStudy (butterflydreamer):

You are basically solving the above ^ Since they've told you that " b = 2" , this means that you just sub (plug) in 2 where the "b" is. Basically REPLACE all the "b's" with 2 and solve :) BUT you want to take note of the absolute value sign also...

OpenStudy (butterflydreamer):

So firstly, let's simplify the parts inside the first absolute sign " -4(2) - 8 " What does -4(2) - 8 = ?

OpenStudy (jessiegonzales):

-8 - 8

OpenStudy (jessiegonzales):

-16?

OpenStudy (butterflydreamer):

yes :) Now can you see how that part was BETWEEN the absolute value sign? Absolute value means you always take the POSITIVE of that value inside. So \[\left| -4(2) - 8 \right| = \left| -8-8 \right|= \left| -16 \right| = 16 \]

OpenStudy (butterflydreamer):

\[\left| -1-(2)^2 \right| = ? \]

OpenStudy (jessiegonzales):

-5

OpenStudy (butterflydreamer):

remember, you take the "absolute value" of -5... So \[\left| -1-(2)^2 \right| = \left| -5 \right| = 5 \] And now solve the last part: \[2(2)^3 = ? \]

OpenStudy (jessiegonzales):

32

OpenStudy (jessiegonzales):

i mean 24

OpenStudy (butterflydreamer):

not quite. What does \[2^3 = ? \]

OpenStudy (jessiegonzales):

16

OpenStudy (butterflydreamer):

err nope.. \[2^3 = 2\times2\times2 = ? \]

OpenStudy (jessiegonzales):

8

OpenStudy (butterflydreamer):

yes and we wanted: \[2(2)^3 = 2\times2\times2\times2 = ?\]

OpenStudy (jessiegonzales):

2(16)

OpenStudy (butterflydreamer):

2(2)^3 = 2 times 2 times 2 times 2 = 16

OpenStudy (jessiegonzales):

2^3=8

OpenStudy (butterflydreamer):

So now we solved the three parts of this equation : \[\left| -4(2) -8 \right| + \left| -1-(2)^2 \right| + 2(2)^3 \] \[= 16 + 5 + 16 = ? \]

OpenStudy (jessiegonzales):

37

OpenStudy (butterflydreamer):

correcttt

OpenStudy (jessiegonzales):

final answer is 37??

OpenStudy (butterflydreamer):

yes

OpenStudy (jessiegonzales):

thankyou zoo much

OpenStudy (butterflydreamer):

you're welcome :)